OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1 = 15 and a15 = -41 ? -157 -153 -149 -145

sam (.sam.):

a15 = -41 a1+(14)d=-41 15+14d=-41 d=-4 a41=a1+(40)d a41=(15)+40(-4)=?

OpenStudy (campbell_st):

use the 15th term to find the common difference \[-41 = 15 +(15 -1)d\] or \[-41 = 15 + 14d\] then the equation for a term in an arithmetic sequence is \[T _{n} = a +(n -1)d\] where n = 41, a = 15 and d was calculated above